Let X

X be N(5,4)

N(5,4) and P(|X−5|>c)=0.01

P(|X−5|>c)=0.01. Then the value of c

c is most approximately equal to


1
Expert's answer
2021-05-07T09:53:49-0400

Given,

X~N(5,4)


Given distribution is a normal distribution with "\mu=5,\sigma=4"

Also, "P(|X-5|>c)=0.01"


"P(|X-5|<c)=1-0.01=0.99"


Hence, The value of c from the table is "2.3"